OpenStudy (anonymous):

[9.01] Identify the vertex for the graph of y = 2x2 + 8x − 3. (2, 21) (2, 17) (−2, −11) (−2, −27) Can someone help me, because I keep getting (-2, -3) and that's obviously not the answer. So can you guys tell me what you're getting?

OpenStudy (anonymous):

Can't you use options?

OpenStudy (anonymous):

I'm not to sure what're you talking about, do you mean the question options?

OpenStudy (anonymous):

yes, you have provided 4 options to us.

OpenStudy (anonymous):

I mean, I have to use the x=-b/2(a) formula to find the x intercept, but every time I plug that into the problem I get something that's not there, so I don't see how I could plug in the options

OpenStudy (anonymous):

you are getting \(x\) as \(-2\) ??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

can't we use this equation: \[y = a(x-h)^2 + k\]

OpenStudy (anonymous):



OpenStudy (anonymous):

This is for axis of symmetry..

OpenStudy (anonymous):

Ive never seen that equation in my class, so I assuming no,I just need help lol, I have been stressing this problem for a while now

OpenStudy (anonymous):

That equation is vertex form of the parabola, there \((h,k)\) is the vertex.

OpenStudy (anonymous):

Okay tell me how you found y = -3??

OpenStudy (anonymous):

I just plugged -2 into the problem, I was taught to do that lol

OpenStudy (anonymous):

Let me check it from the vertex form too.. :)

OpenStudy (anonymous):

Ok :) I hope i'm not coming off snippy or mean, I am really grateful for your help, it's hard to convey emotions over text.

OpenStudy (anonymous):

I found your mistake..

OpenStudy (anonymous):

Plug in \(-2\) in the equation back: \[y = 2(-2)^2 + 8(-2) - 3 \implies y = 8 - 16 - 3 \implies y = -8-3 = ??\]

OpenStudy (anonymous):

Getting?

OpenStudy (anonymous):

@LeilaA7X found it or not??

OpenStudy (anonymous):

Sorry I was doing another problem, lemme work this out

OpenStudy (anonymous):

I got -11

OpenStudy (anonymous):

For more info, I can tell that we can also find vertex from vertex equation.. I will tell you that how to do that.. Yeah \(-11\) is right.. :)

OpenStudy (anonymous):

So the answer is (-2, -11)

OpenStudy (anonymous):

And I understand it for the most part, I just had a slip up on this one lol :)

OpenStudy (anonymous):

Vertex form just need the concept of completing the square..

OpenStudy (anonymous):

Ok, thank you waterineyes!

OpenStudy (anonymous):

\[y = a(x-h)^2 + k \\ y = 2x^2 + 8x - 3 \\ \frac{y}{2} = x^2 + 4x - \frac{3}{2} \\ \frac{y}{2} = x^2 + 4x - \frac{3}{2} + (2)^2 - (2)^2 \\ \frac{y}{2} = (x+2)^2 - \frac{11}{2} \\ \color{blue}{y = 2(x+2)^2 - 11}\]

OpenStudy (anonymous):

It is just like : \(y = a(x-h)^2 + k\), comparison gives you : \((h,k) = (-2, -11)\)

OpenStudy (anonymous):

you are welcome dear.. :)

OpenStudy (anonymous):

I think you will soon learn this too.. :)